# Rapid Word Learning Through Meta In-Context Learning

### v1

@@ Appendix F Evaluation of Generated New Usage Examples
As we mentioned in Section 5.3, we sample study examples from two datasets: the BabyLM-10M test portion and the Chimera dataset (Lazaridou et al., 2017). Statistics of the BabyLM-10M test portion is in Table 5, Appendix A. The Chimera dataset contains 33 chimeras. For each chimera, we merge all examples in its 10 “passages” (each with a different pseudo-word in place of the word) in the original dataset, so we only have 33 different words to learn. We do not use the pseudo-words in the original dataset.
Tables 10 and 11 show samples of example sentences from the BabyLM-10M test portion and the Chimera dataset along with the new usage examples generated by both the Llama-3 8B baseline and the Minnow model finetuned from that baseline. Table 12 contains error analysis of new examples generated from the BabyLM-10M test portion, extending Table 2 in Section 5.3.
@@
$\bullet$ the first blacksmiths were [new-token]. $\bullet$ many civilisations were in the area that is now turkey, like the [new-token], the roman empire and the byzantine empire. $\bullet$ spread of hepatoscopy and astrology to [new-token], etruscans, greeks and romans and to china $\bullet$ the first major empire in the area was the [new-token] (from the 18th century to the 13th century bce). 1. the [new-token] were a people who lived in the area of turkey. 2. perhaps the most famous and widely used alchemical symbol, first popularized by [new-token] alchemists, is the ouroboros. 1. the [new-token] was a major power in the region, and its influence spread to the surrounding areas. 2. epistemology will suffer the [new-token] syndrome. hittites The Minnow model learns that hitties is an ancient ethnic group. However, the greedy-decoded example copies the information (turkey) from the study example, while the sampled example makes seemingly plausible but factually incorrect generalizations (the earliest known ouroboros is found in ancient Egyptian text.) The baseline also generates factual errors (Hittite plague exists but no “hitties syndrome”).
@@ Appendix G Evaluation of Generated Definitions
Additionally, as we also mentioned in Section 5.4, we have additional references of what can be achieved by specialized definition-generation models: the series of FLAN-T5 (Chung et al., 2024) models finetuned by Giulianelli et al. (2023) specifically on generating definitions. This also follows what Teehan et al. (2024) did. These models were finetuned on three corpora, including the Oxford training set (Gadetsky et al., 2018). The series of finetuned FLAN-T5 are listed on their GitHub page (https://github.com/ltgoslo/definition_modeling?tab=readme-ov-file#definition-generation-models-for-english) and can be accessed through Hugging Face model hub. When evaluating the FLAN-T5 models, a pseudo-word ‘wug’ is used as the placeholder for the new word, like in other baselines (Section 3.4) for a fair comparison. Each FLAN-T5 model is prompted with an example sentence followed by a question, “What is the definition of wug?”, as what Giulianelli et al. (2023) did.
Table 13 shows the results of comparing the model-generated and ground-truth definitions from all models, supplementing the brief results in Table 3 with those from the additional specialized FLAN-T5 baselines and the CoLLEGe model. Table 14 shows the average of 1-, 2-, and 3-shot results on the CoLLEGe-DefGen dataset. Tables 15 and 16 show additional definitions generated from the CoLLEGe-DefGen and Oxford test set by the baselines and the Minnow models (in addition to Table 4 in Section 5.4).
Results of CoLLEGe (Teehan et al., 2024), which generates new embeddings to be used in an LLM, are appended to Table 13 and 14 and are directly copied from the original paper. Those numbers should be compared with other models with caution because they have different settings: They are based on Llama-2 7B (Touvron et al., 2023) and new embeddings, their data processing is not as fine as ours (they did not remove duplicated words from both datasets, did not keep the inflectional suffixes in CoLLEGe-DefGen, and did not find more forms in the Oxford dataset as we do (15% examples without replacing any word form)), and the usage examples they randomly selected from CoLLEGe-DefGen are different from ours.
@@
FLAN-T5 Base +DefInstr baseline 83.1 13.1 84.4 16.5
FLAN-T5 Large +DefInstr baseline 83.8 15.5 84.7 17.4
FLAN-T5 XL +DefInstr baseline 83.1 12.4 84.9 19.4
@@
Llama-2 7B CoLLEGe* 84.1 18.0 83.6 17.1
Table 13: Quantitative evaluation of generated definitions by comparing them with ground-truth definitions. This table extends Table 3 by presenting results from all models we evaluate, including the additional specialized FLAN-T5 baselines from Giulianelli et al. (2023) and the CoLLEGe model from Teehan et al. (2024). We generate a definition from only one example (1-shot). We sample an example per word from CoLLEGe-DefGen, while Oxford has exactly one example per word. All definitions are generated with greedy decoding. “+DefInstr” means the definition generation finetuning by Giulianelli et al. (2023). “baseline” means using a pseudo-word ‘wug’ as the placeholder word. For Minnow models (“+Minnow”), scores are averaged across 3 runs. The instruction-tuned variant of Llama-3 8B (“Llama-3 8B Instruct”) is better than the pre-trained variant (“Llama-3 8B”) on definition generation likely due to its better instruction-following ability. *: CoLLEGe results are from “Prompting + CoLLEGe” in the Table 4 of Teehan et al. (2024), which provides Llama-2 7B (Touvron et al., 2023) with embeddings generated by CoLLEGe and prompt it to generate definitions with in-context usage examples. Teehan et al. (2024) has slightly different data processing, so CoLLEGe results are not strictly comparable (see Appendix G).
@@
Llama-2 7B CoLLEGe* 84.8 17.8
Table 14: Quantitative evaluation of generated definitions by comparing them with ground-truth definitions in the CoLLEGe-DefGen dataset. Definitions are generated 1-, 2-, and 3-shot and scores are averaged. All definitions are generated with greedy decoding. For models finetuned with Minnow, scores are averaged across 3 runs. *: CoLLEGe results are from Teehan et al. (2024), which is based on Llama-2 7B and slightly different data processing (see Appendix G). We do not have FLAN-T5 models here since Giulianelli et al. (2023) finetuned them to use only one usage example.
@@
Example Sentence True Definition Minnow Baseline FLAN-XL-DefInstr Word
@@
many people in the world have to [new-token] on \$ 1 a day support oneself to live on something, especially food, in a way that is not very good or healthy to struggle to make ends meet live on subsist
@@
this food [new-token] the patient restore strength to make something more intense or strong a type of food to give strength revitalizes
@@
he is always well-[new-token] care for one’s external appearance well - dressed a type of wug in good health groomed
@@
the lawyers tried to [new-token] the credibility of the witnesses challenge the honesty or veracity of to make something more convincing or believable to question the credibility of a witness to challenge the honesty or veracity of impeach
@@
the car squeaks to a halt and she glares at him because of his [new-token] stop. characterized by abrupt stops and starts a sudden, sharp, high - pitched sound, especially one made by a car’s brakes or a bird’s call a made-up word a jerk that causes an object to move abruptly jerky
@@
try the full plate pork [new-token] : tender pork, oregano-spiked greek salad, warm puffy pita, rice, and aromatic tzatziki-topped lemon potatoes. a greek dish of pieces of meat grilled on a skewer a dish of meat, usually pork, served with a sweet and sour sauce, and often served with rice and vegetables a type of dish that is a combination of pork, rice, and potatoes, typically served with a side of salad and pita bread. a greek dish of grilled meat served in a pita . souvlaki
@@
extend the tv antenna (word is absent) extend or stretch out to a greater or the full length a small, usually round, piece of metal or plastic used to connect two wires together a type of bird raise or extend vertically stretch
@@
the red light gave the central figure increased emphasis (word is absent) special importance or significance a red light a wug is a wug special importance or significance accent
@@
Table 16: Definitions generated by the instruction-tuned variant of Llama-3 8B (baseline), the Minnow model finetuned from it with greedy decoding, and FLAN-XL-DefInstr (i.e., FLAN-T5 XL +DefInstr baseline), using the prompt “The word [new-token] in the above sentence(s) is defined as "” ([new-token] can be replaced by other placeholders, as we mentioned in Section 5.4). Each definition is generated using the single example sentence shown and provided in context. The Minnow model generates reasonable definitions given the context, but is often much longer than the ground-truth definitions, likely because it is not fitted to this dataset. The baseline model is often generating low-quality or repetitive definitions, and sometimes sticks to its prior knowledge of the pseudo-word “wug.” FLAN-XL-DefInstr generates definitions pretty close to the ground-truth, but is sometimes suspicious of overfitting to or memorizing the data, as its definition for ‘impeach’ and ‘accent’ (absent in the example) may suggest.
@@ Appendix H Concepts of “Word”
The term “word” can refer to linguistic units with nuanced variations. Here, we describe the concepts of “word” in different contexts of the paper and their implications. Surprisingly, our models are somehow robust to these variations of “word,” though future work may further improve the processing of words.
@@
In our baselines (Section 3.4 and the FLAN-T5 models in Section 5.4) and comparison of generations (Appendix E), we replace the word to learn by a pseudo-word, like “dax” or “wug”, regardless of the word’s syntactic category and other aspects of meaning. The pseudo-word is then tokenized, usually by a subword tokenizer for LLMs (thus may have multiple tokens). We choose the pseudo-word to be meaningless and commonly used in linguistic tests. However, a pre-trained LLM like Llama-3 may have priors of certain aspects of the pseudo-word’s meaning based on its form. One aspect of the meaning is syntax. For example, from the sentence “Susie goes skiing in the winter”, we replace “skiing” with “dax” and have the sentence “Susie goes dax in the winter.” The sentence has a problem: the part of speech of “skiing” is gerund, but “dax” does not look like a gerund (since it does not end in “-ing”). So the sentence could mislead an LLM like Llama-3, which can use morphological information from its subword tokenization. Another aspect of the meaning is semantics. For example, in Table 16, the baseline model sometimes sticks to its prior knowledge of the pseudo-word “wug,” as reflected in its generated definitions like “a made-up word” and “a type of bird” (“wug” referred to a bird-like creature in the Wug Test of Berko, 1958). We admit that this problem may weaken our baselines and comparison of generations. Future work should use more suitable pseudo-words, preserving the morphological inflections while removing the semantic information.

### v2

@@ Appendix F Evaluation of Generated New Usage Examples
As we mentioned in Section 5.3, we sample study examples from two datasets: the BabyLM-10M test portion and the Chimera dataset (Lazaridou et al., 2017). Statistics of the BabyLM-10M test portion are in Table 5, Appendix A. The Chimera dataset contains 33 chimeras. A chimera is a mixture of two existing and related concepts (e.g., cello and bagpipe). The usage examples of a chimera are sentences using one of the components of the chimera, randomly extracted from a large corpus. For each chimera, we merge all examples in its 10 “passages” (each with a different pseudo-word in place of the word) in the original dataset, so we only have 33 different words to learn. We do not use the pseudo-words in the original dataset.
Tables 11 and 12 show samples of example sentences from the BabyLM-10M test portion and the Chimera dataset along with the new usage examples generated by both the Llama-3 8B baseline and the Minnow model finetuned from that baseline. Table 14 contains error analysis of new examples generated from the BabyLM-10M test portion, extending Table 2 in Section 5.3. For comparing the Minnow model finetuned from Llama-2 7B and the CoLLEGe model on Llama-2 7B, we provide examples generated from the Chimera dataset and the BabyLM-10M test portion in Table 13 and 15, respectively.
@@
Study Example Sentences Minnow CoLLEGe Judgment and Reason Word
$\bullet$ Animals such as capybara, jaguars, jacare ([new-token]s) and hyacinth macaws are particularly vulnerable. $\bullet$ Nadirpur stared at it as though it were a [new-token], his face quite drained. $\bullet$ The owner will be summonsed to court for illegal possession of a [new-token] without a permit. $\bullet$ I went to the cavern to get some, and found a [new-token] in there. $\bullet$ The [new-token] was a big one, and it was a good thing i had a big knife. $\bullet$ 1. The [new-token] is a large, powerful, and dangerous animal. CoLLEGe wins. The Minnow model’s generation contains unrelated information. alligator + rattlesnake
$\bullet$ Fourteen-year-old Craig Ralphs and eighteen year old Glenn Cameron were subsequently arrested and charged with reckless driving and [new-token]s without consent. $\bullet$ " I came in the same [new-token] as someone called Gleeson,"said Marcus. $\bullet$ The chairman of [new-token]s has admitted the company is no longer a mass manufacturer. $\bullet$ She had to continue driving the [new-token] all day, and again the following morning. $\bullet$ The [new-token] was a 1972 ford escort. $\bullet$ 1. The [new-token] is a large, heavy, and powerful bird. Minnow wins. The Minnow model’s generation is very reasonable, while the CoLLEGe model’s generation is incorrect. car + van
$\bullet$ He is principal [new-token] of the English Baroque soloists, with whom he has recorded for Erato and Deutsche Grammophon. $\bullet$ Banners point to the military aspect of [new-token] music: a Jacobite flag and the first Union flag. $\bullet$ " Alexander Baillie said that he was impressed by her playing and even tried out her [new-token]! $\bullet$ Where did the [new-token] idea at the end of Rubina’s Blue Sky Happiness come from? $\bullet$ The [new-token] is a very important part of the music. $\bullet$ 1. The [new-token] of the house was a very old man. Minnow wins. CoLLEGe fails to use the word and generates a default sentence. cello + bagpipe
$\bullet$ 4. Allow the stock to settle then skim off the fat using [new-token] or fat separator. $\bullet$ Add the pasta, cook it for about 12 minutes until just soft, then drain it in a [new-token]. $\bullet$ Imagine you’ve got a bucketful of praise in every room of the house with a big [new-token]. $\bullet$ Plants can be grown in tyres, old metal pots such as buckets, watering cans or even a [new-token]! $\bullet$ I’m going to put a [new-token] on the top of the stairs. $\bullet$ 1. The [new-token] is a small, round, flat, and smooth stone. Tie. The word refers to a kitchen utensil, likely a container, but neither model illustrates this. GPT-4o decides the Minnow as the winner. ladle + colander
$\bullet$ At about the same age I was in a cafe and I spilt some pop on my [new-token]s. $\bullet$ It’s also the first International home [new-token] in history that is reversible as well. $\bullet$ Such as how if you have sufficiently-stretchy pants you could take them off without first removing your [new-token]s. $\bullet$ It was in a cinema on 86th Street where he sat in the balcony, wearing an old sports jacket and open-necked [new-token]. $\bullet$ I’m not sure if it’s a good idea to wear [new-token] in the shower. $\bullet$ 1. The [new-token] is a very comfortable shoe. CoLLEGe wins. Both examples use the word as something wearable, but the thing is clearly not wearable in the shower. trouser + shirt
Table 13: New examples generated on study examples from the Chimera dataset (Lazaridou et al., 2017) by the Minnow model and the CoLLEGe baseline on Llama-2 7B, with greedy decoding. Like in Table 11, the first author provides the judgment and reason (Judgment and Reason).
@@
$\bullet$ the first blacksmiths were [new-token]. $\bullet$ many civilisations were in the area that is now turkey, like the [new-token], the roman empire and the byzantine empire. $\bullet$ spread of hepatoscopy and astrology to [new-token], etruscans, greeks and romans and to china $\bullet$ the first major empire in the area was the [new-token] (from the 18th century to the 13th century bce). 1. the [new-token] were a people who lived in the area of turkey. 2. perhaps the most famous and widely used alchemical symbol, first popularized by [new-token] alchemists, is the ouroboros. 1. the [new-token] was a major power in the region, and its influence spread to the surrounding areas. 2. epistemology will suffer the [new-token] syndrome. hittites The Minnow model learns that hittites is an ancient ethnic group. However, the greedy-decoded example copies the information (turkey) from the study example, while the sampled example makes seemingly plausible but factually incorrect generalizations (the earliest known ouroboros is found in ancient Egyptian text.) The baseline also generates factual errors (Hittite plague exists but no “hittites syndrome”).
@@
Study Example Sentences Minnow CoLLEGe Word Analysis
$\bullet$ the first blacksmiths were [new-token]. $\bullet$ many civilisations were in the area that is now turkey, like the [new-token], the roman empire and the byzantine empire. $\bullet$ spread of hepatoscopy and astrology to [new-token], etruscans, greeks and romans and to china $\bullet$ the first major empire in the area was the [new-token] (from the 18th century to the 13th century bce). 1. the [new-token] were a civilization in the area that is now turkey. 2. the [new-token] empire was destroyed by the medes and persians. 1. The [new-token] were a people of the ancient world. 2. hittites The Minnow model learns that hittites is an ancient ethnic group. However, the greedy-decoded example copies the information (turkey) from the study example, while the sampled example makes seemingly plausible but factually incorrect generalizations. The CoLLEGe baseline’s generation contains even less information, and the sampled result is empty (other sampled results not shown here are in Greek).
$\bullet$ but, i know from his childhood, he says that his father always buys and [new-token] cars just as entertainment, i guess. $\bullet$ in the end, yuji accepts shana’s feelings and the two kiss, which activates a spell left to yuji by crimson lord lamia that [new-token] his existence, so yuji is no longer a torch. $\bullet$ – a medicine which [new-token] health and energy, as extract of malt, cod liver oil, etc. $\bullet$ that diversifies energy supply, [new-token] ecosystems, protects biodiversity, and sequesters carbon. 1. the [new-token] of the bacteria is a result of the action of the enzyme. 2. renewable energy [new-token] supply for energy. 1. The word ’[new-token]’ in example sentences. 2. Sample sentences from outside English. restores For the Minnow model, the greedy-decoded example uses the verb with incorrect part-of-speech but pertinent semantics, but the sampled example is correct and reasonable. The CoLLEGe baseline fails to follow the instructions.
$\bullet$ when i was a student [new-token], i went to the prado in Madrid. $\bullet$ [new-token] is hiking in the outdoors with a backpack full of all your food, clothing, water, and shelter. $\bullet$ [new-token] may be done on short or long trips. $\bullet$ [new-token] (wilderness) 1. [new-token] is a type of outdoor recreation. 2. the song was first performed in 1933 by the hilda [new-token] choral society. 1. The [new-token] of the house was a very old man. 2. 10 pounds of ___________ in ten days. back-packing The Minnow model correctly uses the word in the greedy-decoded example, but fails in the sampled example. The baseline fails to understand the word and generates a default sentence in the greedy-decoded example.
Table 15: Error analysis of new examples generated from the BabyLM-10M test portion by the Minnow model and the CoLLEGe baseline with Llama-2 7B. In each column of generated examples, the first one is generated by greedy decoding, and the second one by sampling with top-p=$0.92$.
@@ Appendix G Evaluation of Generated Definitions
Additionally, as we also mentioned in Section 5.4, we have additional references of what can be achieved by specialized definition-generation models: the series of FLAN-T5 (Chung et al., 2024) models finetuned by Giulianelli et al. (2023) specifically on generating definitions. This also follows what Teehan et al. (2024) did. These models were finetuned on three corpora, including the Oxford training set (Gadetsky et al., 2018). The series of finetuned FLAN-T5 are listed on their GitHub page (https://github.com/ltgoslo/definition_modeling?tab=readme-ov-file#definition-generation-models-for-english) and can be accessed through Hugging Face model hub. When evaluating the FLAN-T5 models, a pseudo-word ‘wug’ is used as the placeholder for the new word, like in other off-the-shelf baselines (Section 3.4.1) for a fair comparison. Each FLAN-T5 model is prompted with an example sentence followed by a question, “What is the definition of wug?”, as what Giulianelli et al. (2023) did.
Table 16 shows the full set of results of comparing the model-generated and ground-truth definitions from all models. Table 17 shows the average of 1-, 2-, and 3-shot results on the CoLLEGe-DefGen dataset. Tables 18 and 19 show additional definitions generated from the CoLLEGe-DefGen and Oxford test set by the baselines and the Minnow models (in addition to Table 4 in Section 5.4).
@@
Llama-2 7B baseline 84.4 14.7 83.9 13.0
+CoLLEGe 84.0 16.3 83.3 14.1
+Minnow 82.9 18.0 83.6 15.6
FLAN-T5 Base +DefInstr baseline 83.1 13.1 84.4 16.5
FLAN-T5 Large +DefInstr baseline 83.8 15.5 84.7 17.4
FLAN-T5 XL +DefInstr baseline 83.1 12.4 84.9 19.4
Table 16: Quantitative evaluation of generated definitions by comparing them with ground-truth definitions. This table extends Table 3 in the main text by adding additional results of the Llama-2 7B baseline and FLAN-T5 models. The FLAN-T5 models generally perform better than all other models on the Oxford dataset, but note that the Oxford dataset is in-distribution for these models, and these models may be overfitting to this dataset (see Table 19 for examples and discussion).
@@
Llama-2 7B baseline 85.2 17.0
+Minnow 84.0 19.9
+CoLLEGe 84.2 16.9
Table 17: Quantitative evaluation of generated definitions by comparing them with ground-truth definitions in the CoLLEGe-DefGen dataset. Definitions are generated 1-, 2-, and 3-shot and scores are averaged. All definitions are generated with greedy decoding. For models finetuned with Minnow, scores are averaged across 3 runs. CoLLEGe* results are from Table 2 of Teehan et al. (2024), which is based on Llama-2 7B and slightly different data processing (see Appendix G). We do not have FLAN-T5 models here since Giulianelli et al. (2023) finetuned them to use only one usage example.
@@
Example Sentence True Definition Minnow Baseline CoLLEGe FLAN-XL-DefInstr Word
@@
many people in the world have to [new-token] on \$ 1 a day support oneself to live on something, especially food, in a way that is not very good or healthy to struggle to make ends meet the [new-token] of two numbers. live on subsist
@@
this food [new-token] the patient restore strength to make something more intense or strong a type of food to make sick or ill. to give strength revitalizes
@@
he is always well-[new-token] care for one’s external appearance well - dressed a type of wug to make a person feel more comfortable or at ease. in good health groomed
@@
the lawyers tried to [new-token] the credibility of the witnesses challenge the honesty or veracity of to make something more convincing or believable to question the credibility of a witness to [new-token] (someone) with a blow or [new-token] (something) by a blow. to challenge the honesty or veracity of impeach
@@
the car squeaks to a halt and she glares at him because of his [new-token] stop. characterized by abrupt stops and starts a sudden, sharp, high - pitched sound, especially one made by a car’s brakes or a bird’s call a made-up word a sudden, [new-token], or [new-token] attack of pain. a jerk that causes an object to move abruptly jerky
@@
try the full plate pork [new-token] : tender pork, oregano-spiked greek salad, warm puffy pita, rice, and aromatic tzatziki-topped lemon potatoes. a greek dish of pieces of meat grilled on a skewer a dish of meat, usually pork, served with a sweet and sour sauce, and often served with rice and vegetables a type of dish that is a combination of pork, rice, and potatoes, typically served with a side of salad and pita bread. a dish of meat, fish, or vegetables cooked in a sauce. a greek dish of grilled meat served in a pita . souvlaki
@@
extend the tv antenna (word is absent) extend or stretch out to a greater or the full length a small, usually round, piece of metal or plastic used to connect two wires together a type of bird to [new-token] or [new-token] (a person) with a weapon. raise or extend vertically stretch
@@
the red light gave the central figure increased emphasis (word is absent) special importance or significance a red light a wug is a wug a sudden, violent, and often uncontrollable attack of fear, dread, or apprehension. special importance or significance accent
@@
Table 19: Definitions generated by the instruction-tuned variant of Llama-3 8B (baseline), the Minnow model finetuned from it with greedy decoding, the CoLLEGe model, and FLAN-XL-DefInstr (i.e., FLAN-T5 XL +DefInstr baseline), using the prompt “The word [new-token] in the above sentence(s) is defined as "” ([new-token] can be replaced by other placeholders, as we mentioned in Section 5.4). Each definition is generated using the single example sentence shown and provided in context. The Minnow model generates reasonable definitions given the context, but is often much longer than the ground-truth definitions, likely because it is not fitted to this dataset. The Llama-3 8B baseline is often generating low-quality or repetitive definitions, and sometimes sticks to its prior knowledge of the pseudo-word “wug.” CoLLEGe often generates definitions that contain the [new-token], or fail to understand the word correctly. FLAN-XL-DefInstr generates definitions pretty close to the ground-truth, but is sometimes suspicious of overfitting to or memorizing the data, as its definition for ‘impeach’ and ‘accent’ (absent in the example) may suggest.
@@ Appendix H Concepts of “Word”
The term “word” can refer to linguistic units with nuanced variations. Here, we describe the concepts of “word” in different contexts of the paper and their implications. Surprisingly, our models are somehow robust to these variations of “word.” Future work may further improve the processing of words and conduct targeted evaluations of morphological variations of the learned words.
@@
In our off-the-shelf baselines (Section 3.4.1 and the additional specialized FLAN-T5 models in Section 5.4) and comparison of generations (Appendix E), we replace the word to learn by a pseudo-word, like “dax” or “wug”, regardless of the word’s syntactic category and other aspects of meaning. The pseudo-word is then tokenized, usually by a subword tokenizer for LLMs (thus may have multiple tokens). We choose the pseudo-word to be meaningless and commonly used in linguistic tests. However, a pre-trained LLM like Llama-3 may have priors of certain aspects of the pseudo-word’s meaning based on its form. One aspect of the meaning is syntax. For example, from the sentence “Susie goes skiing in the winter”, we replace “skiing” with “dax” and have the sentence “Susie goes dax in the winter.” The sentence has a problem: the part of speech of “skiing” is gerund, but “dax” does not look like a gerund (since it does not end in “-ing”). So the sentence could mislead an LLM like Llama-3, which can use morphological information from its subword tokenization. Another aspect of the meaning is semantics. For example, in Table 19, the baseline model sometimes sticks to its prior knowledge of the pseudo-word “wug,” as reflected in its generated definitions like “a made-up word” and “a type of bird” (“wug” referred to a bird-like creature in the Wug Test of Berko, 1958). We admit that this problem may weaken our baselines and comparison of generations. Future work should use more suitable pseudo-words, preserving the morphological inflections while removing the semantic information.
@@
Appendix I Changes in Other Capabilities
How does Minnow finetuning affect other capabilities of language models? As we mentioned in Section 5, we finetune only the input and output embeddings of the two special tokens while freezing all other model parameters. Therefore, we expect that Minnow finetuning will not change other general capabilities of the finetuned language model. To validate this, we evaluate the pre-trained Llama-3 8B and the Minnow finetuned from it on the BLiMP benchmark (Warstadt et al., 2020), which evaluates the grammatical capabilities of language models. Results are shown in Table 20. We find that Minnow finetuning does not change the accuracies very much on most subsets in the benchmark: Most accuracies does not change or change within 0.3%, except for the subset matrix_question_npi_licensor_present, which has a 9.7% decrease and high variance in accuracy. These results reflect that other capabilities of language models are almost unaffected by Minnow.
Phenomenon UID Llama-3 8B +Minnow
anaphor agreement anaphor_gender_agreement 98.9 98.9(0.0)
anaphor_number_agreement 99.5 99.5(0.0)
argument structure animate_subject_passive 80.7 80.7(0.1)
animate_subject_trans 84.2 84.1(0.2)
causative 76.7 76.7(0.0)
drop_argument 79.8 79.8(0.0)
inchoative 70.6 70.6(0.0)
intransitive 83.7 83.7(0.0)
passive_1 90.5 90.5(0.0)
passive_2 90.8 90.8(0.0)
transitive 90.1 90.1(0.0)
binding principle_A_c_command 80.3 80.3(0.0)
principle_A_case_1 100.0 100.0(0.0)
principle_A_case_2 93.8 93.9(0.1)
principle_A_domain_1 99.3 99.3(0.0)
principle_A_domain_2 88.3 88.3(0.0)
principle_A_domain_3 52.8 52.7(0.1)
principle_A_reconstruction 45.3 45.3(0.0)
control/raising existential_there_object_raising 85.0 85.0(0.0)
existential_there_subject_raising 89.9 89.9(0.0)
expletive_it_object_raising 80.3 80.3(0.0)
tough_vs_raising_1 68.8 68.8(0.0)
tough_vs_raising_2 87.2 87.2(0.0)
determiner-noun agreement determiner_noun_agreement_1 99.5 99.5(0.0)
determiner_noun_agreement_2 99.0 99.0(0.0)
determiner_noun_agreement_irregular_1 96.9 96.9(0.0)
determiner_noun_agreement_irregular_2 96.8 96.8(0.0)
determiner_noun_agreement_with_adj_1 97.5 97.5(0.0)
determiner_noun_agreement_with_adj_2 95.4 95.4(0.0)
determiner_noun_agreement_with_adj_irregular_1 92.5 92.5(0.0)
determiner_noun_agreement_with_adj_irregular_2 94.9 94.9(0.0)
ellipsis ellipsis_n_bar_1 79.6 79.6(0.0)
ellipsis_n_bar_2 92.7 92.7(0.1)
filler gap wh_questions_object_gap 81.9 81.9(0.0)
wh_questions_subject_gap 91.7 91.7(0.0)
wh_questions_subject_gap_long_distance 88.0 88.1(0.0)
wh_vs_that_no_gap 97.2 97.2(0.0)
wh_vs_that_no_gap_long_distance 95.6 95.6(0.0)
wh_vs_that_with_gap 39.9 39.8(0.1)
wh_vs_that_with_gap_long_distance 31.7 31.7(0.0)
irregular forms irregular_past_participle_adjectives 95.5 95.5(0.0)
irregular_past_participle_verbs 88.2 88.2(0.0)
island effects adjunct_island 88.5 88.5(0.0)
complex_NP_island 63.2 63.1(0.1)
coordinate_structure_constraint_complex_left_branch 72.0 72.0(0.1)
coordinate_structure_constraint_object_extraction 85.7 85.7(0.0)
left_branch_island_echo_question 42.6 42.5(0.2)
left_branch_island_simple_question 84.8 84.8(0.1)
sentential_subject_island 50.2 50.3(0.0)
wh_island 79.7 79.7(0.0)
npi licensing matrix_question_npi_licensor_present 81.7 72.0(8.7)
npi_present_1 61.6 61.6(0.0)
npi_present_2 70.1 70.1(0.0)
only_npi_licensor_present 97.8 97.5(0.2)
only_npi_scope 88.5 88.2(0.3)
sentential_negation_npi_licensor_present 99.5 99.5(0.0)
sentential_negation_npi_scope 71.4 71.4(0.0)
quantifiers existential_there_quantifiers_1 98.7 98.7(0.0)
existential_there_quantifiers_2 67.1 67.1(0.0)
superlative_quantifiers_1 94.2 94.2(0.0)
superlative_quantifiers_2 90.0 90.2(0.3)
subject-verb agreement distractor_agreement_relational_noun 87.5 87.6(0.0)
distractor_agreement_relative_clause 73.7 73.7(0.1)
irregular_plural_subject_verb_agreement_1 92.1 92.1(0.0)
irregular_plural_subject_verb_agreement_2 94.4 94.4(0.0)
regular_plural_subject_verb_agreement_1 94.3 94.3(0.0)
regular_plural_subject_verb_agreement_2 93.6 93.6(0.0)
NaN Mean 83.5 83.3(0.1)
Table 20: Accuracies on BLiMP (Warstadt et al., 2020). We show the mean and the standard deviation (in the bracket) of 3 runs of Minnow. Minnow accuracies are very close to the pre-trained model.